# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.6.1 Average Flow Approximation Series (AFAS)
$\qquad+2\int_{0}^{t}(t-s)\biggl{[}\biggl{\|}\frac{\operatorname{d}}{\operatorname{d}t}\biggl{(}t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr{)}(s)\biggr{\|}^{2}$
$\qquad+\biggl{<}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl{(}t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr{)}(s)\biggr{>}\biggr{]}\operatorname{d}s.$
@@
$\qquad=\int_{\mathcal{X}^{2,\infty}(M)}p\|x-y\|^{p-1}\bigl{<}x-y,X(x)\bigr{>}\operatorname{d}\tau^{-\overline{V}[\mu]}_{*}V[\mu](X)$
$\qquad-\int_{\mathcal{X}^{2,\infty}(M)}p\|x-y\|^{p-1}\bigl{<}x-y,Y(y)\bigr{>}\operatorname{d}\tau^{-\overline{V}[\nu]}_{*}V[\nu](Y)=0.$
@@
$\qquad+\|\operatorname{d}X(y)(X(y))-\operatorname{d}Y(y)(Y(y))\|$
@@
For the first inequality, we choose a rough estimate of the integral term in (3.68) to get the existence of $C>0$ such that, for $t<1$,
@@
$\qquad+2^{p-2}\int_{0}^{t}(t-s)\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\int_{M\times M}p(p-1)\|x-y\|^{p-2}\|X-Y\|_{W^{2,\infty}(M)}^{2}$
$\qquad+4pR\|x-y\|^{p-1}\|X-Y\|_{W^{2,\infty}(M)}+2pR^{2}\|x-y\|^{p}$
$\qquad+C\max(1,R^{p+2})\biggl{(}\|X-Y\|_{W^{2,\infty}(M)}^{2}+\|X-Y\|_{W^{2,\infty}(M)}+\|x-y\|^{p-2}$
$\qquad+\|x-y\|^{p-1}\biggr{)}t^{p-2}\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s.$
Now, for all $q<p$ and $\mu,\nu\in\mathcal{W}_{p}(M)$, using Proposition 2.4, for an optimal transport plan $g$ from $\mu$ to $\nu$ for the pp–Wasserstein distance, we obtain
$\int_{M\times M}d(x,y)^{p}\operatorname{d}g(x,y)\geq\mathcal{W}_{q}(\mu,\nu)^{p}\geq\biggl{(}\int_{M\times M}d(x,y)^{q}\operatorname{d}g(x,y)\biggr{)}^{\frac{p}{q}}.$ (3.73)
@@
$\qquad\leq\mathcal{W}_{p}(\mu,\nu)^{p}+2^{p-3}p\biggl{(}9(p-1)L^{2}\mathcal{W}_{p}(\mu,\nu)^{p-2}\mathcal{W}_{p}(v_{1},v_{2})^{2}$
[…]
$\qquad+\mathcal{K}\max(1,R^{p+2})\biggl{(}L^{2}\mathcal{W}_{p}(v_{1},v_{2})^{2}+L\mathcal{W}_{p}(v_{1},v_{2})+\mathcal{W}_{p}(\mu,\nu)^{2}+\mathcal{W}_{p}(\mu,\nu)\biggr{)}t^{p}.$
The estimates for $\mathcal{W}_{p}(e^{V[v_{1}]}_{t}(\mu),e^{V[v_{2}]}_{t}(\nu))^{p}$ follows analogously. Now, we work with $p=2$ and give a more precise estimate of the integral term in (3.68). We obtain,
@@
$\qquad=\biggl{<}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\operatorname{d}X(\rho^{X}_{s}(x))\bigl{(}X(\rho^{X}_{s}(x))\bigr{)}-\operatorname{d}X(\rho^{X}_{s}(y))\bigl{(}X(\rho^{X}_{s}(y))\bigr{)}\biggr{>}$
$\qquad+\biggl{<}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\operatorname{d}X(\rho^{X}_{s}(y))\bigl{(}X(\rho^{X}_{s}(y))\bigr{)}-\operatorname{d}Y(\rho^{Y}_{s}(y))\bigl{(}Y(\rho^{Y}_{s}(y))\bigr{)}\biggr{>}$
$\qquad=\biggl{<}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\operatorname{d}X(\rho^{X}_{s}(x))\bigl{(}X(\rho^{X}_{s}(x))\bigr{)}-\operatorname{d}X(\rho^{X}_{s}(y))\bigl{(}X(\rho^{X}_{s}(y))\bigr{)}\biggr{>}$
$\qquad+\biggl{<}x-y+\int_{0}^{s}X\bigl{(}\rho^{X}_{r}(x)\bigr{)}-Y\bigl{(}\rho^{Y}_{r}(y)\bigr{)}\operatorname{d}r,\operatorname{d}X(y)\bigl{(}X(y)\bigr{)}-\operatorname{d}Y(y)\bigl{(}Y(y)\bigr{)}$
$\qquad+\int_{0}^{s}\operatorname{d}^{2}X(\rho^{X}_{r}(y))\bigl{(}X(\rho^{X}_{r}(y))\bigr{)}\bigl{(}X(\rho^{X}_{r}(y))\bigr{)}$
$\qquad+\operatorname{d}X(\rho^{X}_{r}(y))\biggl{(}\operatorname{d}X(\rho^{X}_{r}(y))\bigl{(}\operatorname{d}X(\rho^{X}_{r}(y))\bigr{)}\biggr{)}$
$\qquad-\operatorname{d}^{2}Y(\rho^{Y}_{r}(y))\bigl{(}Y(\rho^{Y}_{r}(y))\bigr{)}\bigl{(}Y(\rho^{Y}_{r}(y))\bigr{)}$
$\qquad+\operatorname{d}Y(\rho^{Y}_{r}(y))\biggl{(}\operatorname{d}Y(\rho^{Y}_{r}(y))\bigl{(}\operatorname{d}Y(\rho^{Y}_{r}(y))\bigr{)}\biggr{)}\operatorname{d}r\biggr{>}.$
@@
$\qquad+2\|X\|^{2}_{W^{2,\infty(M)}}\bigl{\|}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr{\|}^{2}\biggr{)}+2R\|x-y\|\|X-Y\|_{W^{2,\infty}(M)}$
$\qquad+4R^{2}\|X-Y\|_{W^{2,\infty}(M)}s+4R^{3}\|x-y\|s+8R^{4}s^{2}\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s$
@@
$\qquad+R\int_{M\times M}\|x-y\|\operatorname{d}\gamma(x,y)\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\|X-Y\|_{W^{2,\infty}(M)}\operatorname{d}\Gamma(X,Y)\biggr{)}$
$\qquad+4R^{2}\biggl{(}R\int_{M\times M}\|x-y\|\operatorname{d}\gamma(x,y)$
$\qquad+\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\|X-Y\|_{W^{2,\infty}(M)}\operatorname{d}\Gamma(X,Y)\biggr{)}\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}$
$\qquad+\int_{0}^{t}(t-s)\int_{\bigl{(}\mathcal{X}^{2,\infty}(M)\bigr{)}^{2}}\int_{M\times M}8R^{2}\bigl{\|}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr{\|}^{2}\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s$
@@
$\qquad+4R^{2}\biggl{(}R\mathcal{W}_{2}(\mu,\nu)+3\mathcal{W}_{2}(V[v_{1}],V[v_{2}])\biggr{)}\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}$
$\qquad+\int_{0}^{t}(t-s)\int_{\bigl{(}\mathcal{X}^{2,\infty}(M)\bigr{)}^{2}}\int_{M\times M}8R^{2}\bigl{\|}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr{\|}^{2}\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s,$
@@
$\qquad+3LR\mathcal{W}_{2}(\mu,\nu)\mathcal{W}_{2}(v_{1},v_{2})\biggr{)}+4R^{2}\biggl{(}R\mathcal{W}_{2}(\mu,\nu)+3L\mathcal{W}_{2}(v_{1},v_{2})\biggr{)}\frac{t^{3}}{6}$
$\qquad+\frac{2R^{4}}{3}t^{4}\biggr{)}e^{8R^{2}t^{2}}.$
@@
Assume $(M,g)=(\mathbb{R}^{n},g)$. Let $R>0$ and $v,w\in\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ supported in the centered ball with radius $R$. Then, for all $\mu,\nu$, the estimates (3.56), (3.57), (3.60) and (3.61) also hold.
@@
Let $T>0$, $p\geq 2$, $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\mathcal{}\mathfrak{P}$ be a partition of $[0,T]$. Let $V$ be a vector field probability. Assume that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$. Then, the Average Flow Approximation Series $\mu^{\mathcal{}\mathfrak{P}}$ with initial value $\mu_{0}$ is locally Hölder continuous with respect to the pp–Wasserstein distance with exponent $\frac{1}{p}$. Furthermore, the function $t\mapsto\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})$ is Lipschitz with respect to the pp–Wasserstein distance. More precisely, there exists $\delta>0$ such that, if the path–length $\mathfrak{p}(\mathcal{}\mathfrak{P})\leq\delta$, for all $t,s\in[0,T]$, with $|t-s|\leq\frac{1}{2^{p-4}p\bigl{(}(9p+5)R^{2}+3\bigr{)}}$, we have
$\mathcal{W}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t},\mu^{\mathcal{}\mathfrak{P}}_{s})\leq C(M)\biggl{(}2^{p-4}p\bigl{(}(9p+5)R^{2}+3\bigr{)}\biggr{)}^{\frac{1}{p}}|t-s|^{\frac{1}{p}},$ (3.77)
$\bigl{|}\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})-\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{s})\bigr{|}\leq C(M)2^{p-4}p\bigl{(}(9p+5)R^{2}+3\bigr{)}|t-s|,$

### v2

@@ 3.6.3 Averaged flow operators—Lipschitz measure–dependence
$\hskip 18.49988pt+2\int_{0}^{t}(t-s)\biggl[\biggl\|\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr\|^{2}$
$\hskip 18.49988pt+\biggl<\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr>\biggr]\operatorname{d}s.$
@@
$\hskip 18.49988pt=\int_{\mathcal{X}^{2,\infty}(M)}p\|x-y\|^{p-1}\bigl<x-y,X(x)\bigr>\operatorname{d}\tau^{-\overline{V}[\mu]}_{*}V[\mu](X)$
$\hskip 18.49988pt-\int_{\mathcal{X}^{2,\infty}(M)}p\|x-y\|^{p-1}\bigl<x-y,Y(y)\bigr>\operatorname{d}\tau^{-\overline{V}[\nu]}_{*}V[\nu](Y)=0.$
@@
$\qquad\hskip 18.49988pt+\|\operatorname{d}X(y)(X(y))-\operatorname{d}Y(y)(Y(y))\|$
@@
Combining (3.66) with the estimates
$\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\|\leq\|x-y\|+s\bigl(\|X\|_{W^{2,\infty}(M)}+\|Y\|_{W^{2,\infty}(M)}\bigr),$ (3.71)
$\biggl\|\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr\|\leq\|X\|_{W^{2,\infty}(M)}\|x-y\|+\|X-Y\|_{W^{2,\infty}(M)}$
$\hskip 18.49988pt+s\bigl(\|X\|_{W^{2,\infty}(M)}^{2}+\|Y\|_{W^{2,\infty}(M)}^{2}\bigr),$
$\biggl\|\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr\|\leq 2\|X\|_{W^{2,\infty}(M)}^{2}\|x-y\|$
$\hskip 18.49988pt+2\bigl(\|X\|_{W^{2,\infty}(M)}+\|Y\|_{W^{2,\infty}(M)}\bigr)\|X-Y\|_{W^{2,\infty}(M)}$
$\hskip 18.49988pt+2s\bigl(\|X\|_{W^{2,\infty}(M)}^{3}+\|Y\|_{W^{2,\infty}(M)}^{3}\bigr),$
and subsequently using the inequality of Cauchy–Schwarz, we roughly estimate the integral term in (3.67) to obtain the existence of $C>0$ such that, for $t<1$,
@@
$\hskip 18.49988pt+2^{p-1}\int_{0}^{t}(t-s)\int_{\bigl(\mathcal{X}^{2,\infty}(M)\bigr)^{2}}\int_{M\times M}p(p-1)\|x-y\|^{p-2}\|X-Y\|_{W^{2,\infty}(M)}^{2}$
$\hskip 18.49988pt+2pR\|x-y\|^{p-1}\|X-Y\|_{W^{2,\infty}(M)}+p^{2}R^{2}\|x-y\|^{p}$
$\hskip 18.49988pt+C\max(1,R^{p+2})\biggl(s^{p}+\|X-Y\|_{W^{2,\infty}(M)}^{2}s^{p-2}+\|X-Y\|_{W^{2,\infty}(M)}s^{p-1}$
$\hskip 18.49988pt+\|x-y\|^{p-1}s+\|x-y\|^{p-2}s^{2}+\|x-y\|^{2}s^{p-2}$
$\hskip 18.49988pt+\|x-y\|s^{p-1}+\|x-y\|\|X-Y\|_{W^{2,\infty}(M)}s^{p-2}$
$\hskip 18.49988pt+\|x-y\|^{p-2}\|X-Y\|_{W^{2,\infty}(M)}s\biggr)\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s.$
Now, for all $q<p$ and $\mu,\nu\in\mathcal{W}_{p}(M)$, by Jensen’s inequality, we obtain for an optimal transport plan $g$ from $\mu$ to $\nu$ for the pp–Wasserstein distance,
$\mathcal{W}_{p}(\mu,\nu)^{p}=\int_{M\times M}d(x,y)^{p}\operatorname{d}g(x,y)\geq\biggl(\int_{M\times M}d(x,y)^{q}\operatorname{d}g(x,y)\biggr)^{\frac{p}{q}}.$ (3.73)
@@
$\hskip 18.49988pt\leq\mathcal{W}_{p}(\mu,\nu)^{p}+2^{p-2}p\biggl(9(p-1)\mathcal{W}_{p}(\mu,\nu)^{p-2}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])^{2}$
$\hskip 18.49988pt+pR^{2}\mathcal{W}_{p}(\mu,\nu)^{p}+6R\mathcal{W}_{p}(\mu,\nu)^{p-1}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])\biggr)t^{2}$
$\hskip 18.49988pt+\mathcal{K}\max(1,R^{p+2})\biggl(t^{p+2}+\mathcal{W}_{p}(V[v_{1}],V[v_{2}])^{2}t^{p}+\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{p+1}$
$\hskip 18.49988pt+\mathcal{W}_{p}(\mu,\nu)^{p-1}t^{3}+\mathcal{W}_{p}(\mu,\nu)^{p-2}t^{4}+\mathcal{W}_{p}(\mu,\nu)^{2}t^{p}+\mathcal{W}_{p}(\mu,\nu)t^{p+1}$
$\hskip 18.49988pt+\mathcal{W}_{p}(\mu,\nu)\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{p}++\mathcal{W}_{p}(\mu,\nu)^{p-2}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{3}\biggr).$
In particular, if $V$ is LL–Lipschitz with respect to the $\mathcal{W}_{p}$–distance,
$\mathcal{W}_{p}(f^{V}_{t^{2}}(\mu),f^{V}_{t^{2}}(\nu))^{p}$ (3.75)
$\hskip 18.49988pt\leq\mathcal{W}_{p}(\mu,\nu)^{p}+2^{p-2}p\biggl(9(p-1)L^{2}+pR^{2}+6RL\biggr)\mathcal{W}_{p}(\mu,\nu)^{p}t^{2}$
$\hskip 18.49988pt+\mathcal{K}\max(1,R^{p+2})\biggl(1+\mathcal{W}_{p}(\mu,\nu)^{p-1}+\mathcal{W}_{p}(\mu,\nu)^{p-2}$
$\hskip 18.49988pt+(L^{2}+L+1)\mathcal{W}_{p}(\mu,\nu)^{2}+(1+L)\mathcal{W}_{p}(\mu,\nu)\biggr)t^{\min(3,p)}.$
The estimates for $\mathcal{W}_{p}(e^{V[v_{1}]}_{t}(\mu),e^{V[v_{2}]}_{t}(\nu))^{p}$ follow analogously. Now, we work with $p=2$ and give a more precise estimate of the integral term in (3.67). We get,
@@
$\hskip 18.49988pt=\biggl<\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\operatorname{d}X(\rho^{X}_{s}(x))\bigl(X(\rho^{X}_{s}(x))\bigr)-\operatorname{d}X(\rho^{X}_{s}(y))\bigl(X(\rho^{X}_{s}(y))\bigr)\biggr$>
$\hskip 18.49988pt+\biggl<\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\operatorname{d}X(\rho^{X}_{s}(y))\bigl(X(\rho^{X}_{s}(y))\bigr)-\operatorname{d}Y(\rho^{Y}_{s}(y))\bigl(Y(\rho^{Y}_{s}(y))\bigr)\biggr$>
$\hskip 18.49988pt=\biggl<\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\operatorname{d}X(\rho^{X}_{s}(x))\bigl(X(\rho^{X}_{s}(x))\bigr)-\operatorname{d}X(\rho^{X}_{s}(y))\bigl(X(\rho^{X}_{s}(y))\bigr)\biggr$>
$\hskip 18.49988pt+\biggl<x-y+\int_{0}^{s}X\bigl(\rho^{X}_{r}(x)\bigr)-Y\bigl(\rho^{Y}_{r}(y)\bigr)\operatorname{d}r,\operatorname{d}X(y)\bigl(X(y)\bigr)-\operatorname{d}Y(y)\bigl(Y(y)\bigr)$
$\hskip 18.49988pt+\int_{0}^{s}\operatorname{d}^{2}X(\rho^{X}_{r}(y))\bigl(X(\rho^{X}_{r}(y))\bigr)\bigl(X(\rho^{X}_{r}(y))\bigr)$
$\hskip 18.49988pt+\operatorname{d}X(\rho^{X}_{r}(y))\biggl(\operatorname{d}X(\rho^{X}_{r}(y))\bigl(\operatorname{d}X(\rho^{X}_{r}(y))\bigr)\biggr)$
$\hskip 18.49988pt-\operatorname{d}^{2}Y(\rho^{Y}_{r}(y))\bigl(Y(\rho^{Y}_{r}(y))\bigr)\bigl(Y(\rho^{Y}_{r}(y))\bigr)$
$\hskip 18.49988pt+\operatorname{d}Y(\rho^{Y}_{r}(y))\biggl(\operatorname{d}Y(\rho^{Y}_{r}(y))\bigl(\operatorname{d}Y(\rho^{Y}_{r}(y))\bigr)\biggr)\operatorname{d}r\biggr$>.
@@
$\hskip 18.49988pt+2\|X\|^{2}_{W^{2,\infty(M)}}\bigl\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr\|^{2}\biggr)+2R\|x-y\|\|X-Y\|_{W^{2,\infty}(M)}$
$\hskip 18.49988pt+4R^{2}\|X-Y\|_{W^{2,\infty}(M)}s+4R^{3}\|x-y\|s+8R^{4}s^{2}\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s$
@@
$\hskip 18.49988pt+R\int_{M\times M}\|x-y\|\operatorname{d}\gamma(x,y)\int_{\bigl(\mathcal{X}^{2,\infty}(M)\bigr)^{2}}\|X-Y\|_{W^{2,\infty}(M)}\operatorname{d}\Gamma(X,Y)\biggr)$
$\hskip 18.49988pt+4R^{2}\biggl(R\int_{M\times M}\|x-y\|\operatorname{d}\gamma(x,y)$
$\hskip 18.49988pt+\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\|X-Y\|_{W^{2,\infty}(M)}\operatorname{d}\Gamma(X,Y)\biggr)\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}$
$\hskip 18.49988pt+\int_{0}^{t}(t-s)\!\!\int_{\bigl(\mathcal{X}^{2,\infty}(M)\bigr)^{2}}\!\!\int_{M^{2}}\!\!8R^{2}\bigl\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr\|^{2}\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s$
@@
$\hskip 18.49988pt+4R^{2}\biggl(R\mathcal{W}_{2}(\mu,\nu)+3\mathcal{W}_{2}(V[v_{1}],V[v_{2}])\biggr)\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}$
$\hskip 18.49988pt+\int_{0}^{t}(t-s)\!\!\int_{\bigl(\mathcal{X}^{2,\infty}(M)\bigr)^{2}}\!\!\int_{M^{2}}\!\!8R^{2}\bigl\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr\|^{2}\operatorname{d}\gamma(x,y)\operatorname{d}\Gamma(X,Y)\operatorname{d}s,$
@@
$\hskip 18.49988pt+3LR\mathcal{W}_{2}(\mu,\nu)\mathcal{W}_{2}(v_{1},v_{2})\biggr)+4R^{2}\biggl(R\mathcal{W}_{2}(\mu,\nu)+3L\mathcal{W}_{2}(v_{1},v_{2})\biggr)\frac{t^{3}}{6}$
$\hskip 18.49988pt+\frac{2R^{4}}{3}t^{4}\biggr)e^{8R^{2}t^{2}}.$
@@
An analogon of Lemma 3.33 also holds on all Riemannian manifolds with bounded geometry.
Lemma 3.34.
Let $(M,g)$ be a Riemannian manifold with bounded geometry. With the assumptions of Lemma 3.33, the estimates (3.55), (3.56), (3.59) and (3.60) also hold, with an additional multiplicative constant $C(M)>0$ depending only on the manifold $M$, with $C(\mathbb{R}^{n})=1$.
Proof.
We proceed similarly as in the proof of Lemma 3.33 in local charts. Let $x,y\in M$ and $X,Y\in\mathcal{X}^{2,\infty}(M)$ and $t>0$. If there exists a unique smooth unit–speed geodesic $\gamma$ from $\rho^{X}_{t}(x)$ to $\rho^{Y}_{t}(y)$, a technical computation yields
$\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto d\bigl(\rho^{X}_{t}(x),\rho^{Y}_{t}(y)\bigr)^{p}\biggr)(s)$ (3.78)
$\qquad=pd\bigl(\rho^{X}_{t}(x),\rho^{Y}_{t}(y)\bigr)^{p-1}\biggl[g(\rho^{X}_{t}(x))\biggl(X(\rho^{X}_{t}(x)),\frac{\operatorname{d}}{\operatorname{d}t}\gamma(0)\biggr)$
$\hskip 18.49988pt-g(\rho^{Y}_{t}(y))\biggl(Y(\rho^{Y}_{t}(y)),\frac{\operatorname{d}}{\operatorname{d}t}\gamma\bigl(d(\rho^{X}_{t}(x),\rho^{Y}_{t}(y))\bigr)\biggr)\biggr],$
such that, in particular, if there exists a unique minimizing geodesic between $x$ and $y$,
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl(t\mapsto d\bigl(\rho^{X}_{t}(x),\rho^{Y}_{t}(y)\bigr)^{p}\biggr)(0)$ (3.79)
$\qquad=p(p-2)d\bigl(x,y\bigr)^{p-2}\biggl[g(x)\biggl(X(x),\frac{\operatorname{d}}{\operatorname{d}t}\gamma(0)\biggr)$
$\hskip 18.49988pt\hskip 18.49988pt-g(y)\biggl(Y(y),\frac{\operatorname{d}}{\operatorname{d}t}\gamma(d(x,y))\biggr)\biggr]^{2}$
$\qquad+pd\bigl(x,y\bigr)^{p-2}d^{2}(x,y)\biggl((x,y)\mapsto\frac{1}{2}d(x,y)^{2}\biggr)\biggl[X(x),Y(y),X(x),Y(y)\biggr]$
$\qquad-pd\bigl(x,y\bigr)^{p-2}\biggl[g(x)\biggl(\nabla_{X}X(x),\frac{\operatorname{d}}{\operatorname{d}t}\gamma(0)\biggr)$
$\hskip 18.49988pt\hskip 18.49988pt-g(y)\biggl(\nabla_{Y}Y(y),\frac{\operatorname{d}}{\operatorname{d}t}\gamma(d(x,y))\biggr)\biggr]^{2}.$
Using the bounded geometry of $M$ to unifomly bound the curvature terms appearing in the second derivative above, we directly obtain a similar estimate as the expression inside the integral in (3.72), if $M$ is uniquely geodesic.
If there are multiple minimizing geodesics between $x$ and $y$, the Hessian of $d^{2}:(z,w)\mapsto d(z,w)^{2}$ might not be well–defined in $(x,y)$. However, $d^{2}$ is still $W^{1,\infty}$ on any bounded tubular neighbourhood of the diagonal of $M\times M$. Thus, its Clarke generalised Hessian (see Rockafellar and Wets [26]) is well–defined and uniformly bounded on such a neighbourhood, again by the bounded geometry of $M$. A slightly more involved computation then yields an analogous estimate as the one we obtain for uniquely geodesic manifolds. $\square$
Lemma 3.33 and Lemma 3.34 are phrased for probability vector–fields for convenience. We obtain a similar result for average flow–operators associated to any vector–field probabilities in a completely analogous way.
@@
Let $(M,g)$ be of bounded geometry. Let $v,w\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $R>0$ with $\mathcal{M}_{p+2}(\mu),\mathcal{M}_{p+2}(\nu)\leq R^{p+2}$. Then, for all $\mu,\nu$, the estimates (3.55), (3.56), (3.59) and (3.60) also hold.
@@
Let $T>0$, $p\geq 2$, $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\mathcal{}\mathfrak{P}$ be a partition of $[0,T]$. Let $V$ be a probability vector–field. Assume that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathbb{R}^{n})$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$. Then:
  • The Average Flow Approximation Series $\mu^{\mathcal{}\mathfrak{P}}$ with initial value $\mu_{0}$ is locally Hölder continuous with respect to the pp–Wasserstein distance with exponent $\frac{1}{p}$.
  • More precisely, there exists $\delta>0$ such that, if the path–length $\mathfrak{p}(\mathcal{}\mathfrak{P})\leq\delta$, for all $t,s\in[0,T]$, with $|t-s|\leq\frac{1}{2^{p-3}pR\bigl(5R(2p+1)+4\bigr)}$, we have
    $\mathcal{W}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t},\mu^{\mathcal{}\mathfrak{P}}_{s})\leq C(M)\biggl(2^{p-3}pR\bigl(5R(2p+1)+4\bigr)\biggr)^{\frac{1}{p}}|t-s|^{\frac{1}{p}},$ (3.80)